OpenStudy (anonymous):

What is the value of X? 6/x = 24/28.8

563blackghost (563blackghost):

How can you get from 6 to 24?

OpenStudy (anonymous):

its a fraction they have to be equivalent @563blackghost

563blackghost (563blackghost):

\[\frac{ 24 }{ 28.8 }\div \frac{ ? }{ ? }=\frac{ 6 }{ x }\]

563blackghost (563blackghost):

so.... how do we get from 24 down to 6?

OpenStudy (anonymous):

that's not it and how did you do it like that? @563blackghost

563blackghost (563blackghost):

\[24\div6=?\]

563blackghost (563blackghost):

plz follow thru...

OpenStudy (aripotta):

blackghost is trying to use equivalent ratios to solve the problem. just follow his/her guidance and you should get your answer

OpenStudy (anonymous):

6